OpenStudy (anonymous):

Solve for x: 2x – 6 +x = 13 – 4x LAST ONE, PLEASE SHOW ME HOW TO DO

OpenStudy (anonymous):

3x -6 = 13 - 4x 7x-6=13 7x=19 x=19/7

OpenStudy (anonymous):

THANK YOU SO MUCH!! now i can go to bed lol(: @bearcat14